# On the Convergence of Muon and Beyond

### v2

@@ 7 Conclusion
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t}^{2}L^{2}n+2\eta_{t+1}^{2}\sigma^{2}$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})$
@@
$\sqrt{\eta_{t}}A_{t}\leq 4\left(\frac{A_{t}}{\sqrt{\eta_{t}}}-\frac{A_{t+1}}{\sqrt{\eta_{t+1}}}+\frac{B_{t+1}}{\sqrt{\eta_{t+1}}}\right)$
@@
$\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq P_{t}-P_{t+1}+\frac{4\cdot 2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})}{\sqrt{\eta_{t+1}}}=P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2})$
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}4\sigma^{2}$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
@@
$\langle\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+2Ln+4(2Ln+\sigma^{2}L^{-1})(1+\ln T)+\frac{Ln}{2}(1+\ln T)$
@@
$G=f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+\left(8Ln+4\sigma^{2}L^{-1}+\frac{Ln}{2}\right)(1+\ln T)+2Ln$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}$
@@
$=\mathcal{O}\left(\frac{\ln T}{T^{1/3}}\right)$
@@
Appendix E Lemmas for Theorems 4.3 and 4.4
@@
Let $\{\Gamma_{t}\}_{t\geq 1}$ be a non-negative sequence whose partial sum $D_{t}=\sum_{t=1}^{T}\Gamma_{t}$ is bounded by $D_{t}\leq A_{1}\ln T+A_{2}$ for some positive constants $A_{1}$ and $A_{2}$. Define the function $f(t)=\frac{t^{1-p}}{\ln t}$ where $p\in(0,1)$. Then, for any sufficiently large integer $T_{0}>e^{1/(1-p)}$ where $e$ is the base of the natural logarithm, the following inequality holds for all $T>T_{0}$:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq\left(A_{1}+A_{2}\right)(T-1)^{1-p}$
@@
We choose an integer $T_{0}>e^{1/(1-p)}$ large enough such that for all $t\geq T_{0}$, the function $f(t)$ is positive and monotonically increasing. We use summation by parts to bound the sum $\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)$:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)=\sum_{t=T_{0}}^{T-1}(D_{t}-D_{t-1})f(t)=D_{T-1}f(T-1)-D_{T_{0}-1}f(T_{0})-\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1))$
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}(A_{1}+A_{2})(T-1)^{1-p}$
@@
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}$
where $p\in(0,1)$ and $\mu>0$ are constants. The non-negative noise sequence $\{\Gamma_{t}\}_{t\geq 1}$ satisfies the condition from Lemma E.1, with its partial sum bounded by $\sum_{t=1}^{T}\Gamma_{t}=\mathcal{O}(\ln T)$.
Then, the sequence $\Delta_{t}$ converges to 0 with the following rate:
$\Delta_{t}=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right)$
@@
We first define the variable $G_{t}=\sqrt{\Delta_{t}}$. Since $\Delta_{t}>0$, it follows that $G_{t}>0$. The function $f(x)=\sqrt{x}$ is concave, which implies $\sqrt{y}-\sqrt{x}\leq\frac{1}{2\sqrt{x}}(y-x)$ for any $x,y>0$. By setting $y=\Delta_{t+1}$ and $x=\Delta_{t}$, we obtain:
$G_{t+1}-G_{t}\leq\frac{\Delta_{t+1}-\Delta_{t}}{2G_{t}}$
@@
$G_{t+1}-G_{t}\leq\frac{1}{2G_{t}}\left(-\frac{\sqrt{2\mu}}{t^{p}}G_{t}+\Gamma_{t}\right)$
Letting $C=\frac{\sqrt{2\mu}}{2}$, we arrive at the core recurrence for $G_{t}$:
@@
For $p\in(0,1)$, we will prove that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$ by contradiction.
@@
$G_{T}-G_{T_{0}}\leq\sum_{t=T_{0}}^{T-1}(G_{t+1}-G_{t})\leq\sum_{t=T_{0}}^{T-1}\left(-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}\right)$
@@
$G_{T}<G_{T_{0}}-\underbrace{C\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}}_{V_{1}}+\underbrace{\frac{1}{2A}\sum_{t=T_{0}}^{T-1}\frac{\Gamma_{t}t^{1-p}}{\ln t}}_{V_{2}}$
@@
$G_{T}<G_{T_{0}}-\frac{C}{1-p}T^{1-p}+\frac{\mathcal{A}}{2A}T^{1-p}=G_{T_{0}}+T^{1-p}\left(\frac{\mathcal{A}}{2A}-\frac{C}{1-p}\right)$
@@
$\frac{\mathcal{A}}{2A}-\frac{C}{1-p}<0\implies A>\frac{\mathcal{A}}{2C}$
@@
This contradiction proves that our assumption that $G_{t}$ can remain above $B_{t}$ for an arbitrarily long interval is false. Therefore, there must exist a time $T_{A}$ such that for all $t>T_{A}>T_{0}$, $G_{t}\leq B_{t}$. This establishes that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$.
Finally, squaring both sides of this result yields the convergence rate for $\Delta_{t}$:
$\Delta_{t}=G_{t}^{2}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1-p}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right)$
@@
$=\Delta_{t}-\frac{\sqrt{2\mu}}{t^{3/4}}\sqrt{\Delta_{t}}+\Gamma_{t}$
By Lemma E.2, we have
$G_{t+1}\leq G_{t}-\frac{C}{t^{3/4}}+\frac{\Gamma_{t}}{2G_{t}}$
Setting $p=3/4$, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{8C}$, there exists a time step $T_{A}$ such that for all $t>T_{A}$, we have $G_{t}\leq A\cdot\frac{\ln t}{t^{1/4}}$. Thus, we obtain the convergence rate for $G_{t}$:
$G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1/4}}\right)$
Finally, since $\Delta_{t}=G_{t}^{2}$, squaring both sides yields the final convergence rate of $\Delta_{t}$:
$\Delta_{t}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1/4}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{1/2}}\right)$
Hence,
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right)$
@@
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
@@
$+\underbrace{\frac{L\eta_{t}^{3/2}}{2}+\left(\frac{\epsilon}{2}+\frac{\sqrt{\eta_{t}}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}}_{\Gamma_{t}}$
@@
$\Gamma_{t}=\frac{\sqrt{\eta_{t}}}{L}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+(L/2+Ln)\eta_{t}^{3/2}$
@@
$\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}$
@@
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{2/3}}\sqrt{\Delta_{t}}+\Gamma_{t}$
This recurrence is in the form required by Lemma E.2. Let $G_{t}=\sqrt{\Delta_{t}}$, $C=\sqrt{2\mu}/2$ and $p=2/3$. According to Lemma E.2, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{6C}$, there exists a time $T_{A}$ such that for all $t>T_{A}$, $G_{t}\leq A\cdot\frac{\ln t}{t^{1/3}}$. This establishes the convergence rate for $G_{t}$:
$G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1/3}}\right)$
Finally, since $\Delta_{t}=G_{t}^{2}$, squaring both sides gives the final convergence rate for $\Delta_{t}$:
$\Delta_{t}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1/3}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2/3}}\right)$
Therefore, we have
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right)$

### v3

@@ 6 Conclusion
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t}^{2}L^{2}n+2\eta_{t+1}^{2}\sigma^{2},$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2}).$
@@
$\sqrt{\eta_{t}}A_{t}\leq 4\left(\frac{A_{t}}{\sqrt{\eta_{t}}}-\frac{A_{t+1}}{\sqrt{\eta_{t+1}}}+\frac{B_{t+1}}{\sqrt{\eta_{t+1}}}\right).$
@@
$\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq P_{t}-P_{t+1}+\frac{4\cdot 2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})}{\sqrt{\eta_{t+1}}}=P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2}).$
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T),$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}4\sigma^{2}.$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2},$
@@
$\langle\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F},$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+2Ln+4(2Ln+\sigma^{2}L^{-1})(1+\ln T)+\frac{Ln}{2}(1+\ln T),$
@@
$G=f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+\left(8Ln+4\sigma^{2}L^{-1}+\frac{Ln}{2}\right)(1+\ln T)+2Ln.$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}.$
@@
$=\mathcal{O}\left(\frac{\ln T}{T^{1/3}}\right).$
@@
Appendix E Lemmas for Theorems 3.3 and 3.4
@@
Let $\{\Gamma_{t}\}_{t\geq 1}$ be a non-negative sequence whose partial sum $D_{t}=\sum_{t=1}^{T}\Gamma_{t}$ is bounded by $D_{t}\leq A_{1}\ln T+A_{2}$ for some positive constants $A_{1}$ and $A_{2}$. Define the function $f(t)=\frac{t^{1-p}}{\ln t}$ where $p\in(0,1)$. Fix $T_{0}=2e^{1/(1-p)}$, where $e$ is the base of the natural logarithm. Then, the following inequality holds for all $T>T_{0}$:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq\left(A_{1}+A_{2}\right)(T-1)^{1-p}.$
@@
We choose an integer $T_{0}=2e^{1/(1-p)}>e^{1/(1-p)}$ large enough such that for all $t\geq T_{0}$, the function $f(t)$ is positive and monotonically increasing. We use summation by parts to bound the sum $\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)$:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)=\sum_{t=T_{0}}^{T-1}(D_{t}-D_{t-1})f(t)=D_{T-1}f(T-1)-D_{T_{0}-1}f(T_{0})-\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1)).$
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}(A_{1}+A_{2})(T-1)^{1-p},$
@@
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t},$
where $p\in(0,1)$ and $\mu>0$ are constants. The non-negative noise sequence $\{\Gamma_{t}\}_{t\geq 1}$ satisfies the condition from Lemma E.1 that $\sum_{t=1}^{T}\Gamma_{t}=A_{1}\ln T+A_{2}$. Then, the sequence $\Delta_{t}$ converges to 0 with the following rate:
$\Delta_{t}\leq\frac{\mathcal{A}^{2}(1-p)^{2}}{C^{2}}\cdot\frac{(\ln t)^{2}}{t^{2(1-p)}},$
where $\mathcal{A}:=A_{1}+A_{2}$ and $C=\frac{\sqrt{2\mu}}{2}$.
@@
We define the variable $G_{t}=\sqrt{\Delta_{t}}$.
Since $\Delta_{t}>0$, it follows that $G_{t}>0$. The function $f(x)=\sqrt{x}$ is concave, which implies $\sqrt{y}-\sqrt{x}\leq\frac{1}{2\sqrt{x}}(y-x)$ for any $x,y>0$. By setting $y=\Delta_{t+1}$ and $x=\Delta_{t}$, we obtain:
$G_{t+1}-G_{t}\leq\frac{\Delta_{t+1}-\Delta_{t}}{2G_{t}}.$
@@
$G_{t+1}-G_{t}\leq\frac{1}{2G_{t}}\left(-\frac{\sqrt{2\mu}}{t^{p}}G_{t}+\Gamma_{t}\right).$
Letting $C=\frac{\sqrt{2\mu}}{2},T_{0}=2e^{1/(1-p)}$, we arrive at the core recurrence for $G_{t}$:
@@
For $p\in(0,1)$, we will prove that $G_{t}\leq\frac{\mathcal{A}(1-p)}{C}\cdot\frac{\ln t}{t^{1-p}}$ by contradiction.
@@
$G_{T}-G_{T_{0}}\leq\sum_{t=T_{0}}^{T-1}(G_{t+1}-G_{t})\leq\sum_{t=T_{0}}^{T-1}\left(-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}\right).$
@@
$G_{T}<G_{T_{0}}-\underbrace{C\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}}_{V_{1}}+\underbrace{\frac{1}{2A}\sum_{t=T_{0}}^{T-1}\frac{\Gamma_{t}t^{1-p}}{\ln t}}_{V_{2}}.$
@@
$G_{T}<G_{T_{0}}-\frac{C}{1-p}T^{1-p}+\frac{\mathcal{A}}{2A}T^{1-p}=G_{T_{0}}+T^{1-p}\left(\frac{\mathcal{A}}{2A}-\frac{C}{1-p}\right).$
@@
$\frac{\mathcal{A}}{2A}-\frac{C}{1-p}<0\implies A>\frac{\mathcal{A}(1-p)}{2C}.$
@@
This contradiction proves that our assumption that $G_{t}$ can remain above $B_{t}$ for an arbitrarily long interval is false. Therefore, there must exist a time $T_{A}$ such that for all $t>T_{A}>T_{0}=2e^{1/(1-p)}$, $G_{t}\leq B_{t}$. This establishes that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$.
Finally, by choosing the constant $A=\frac{\mathcal{A}(1-p)}{C}$ , squaring both sides of this result yields the convergence rate for $\Delta_{t}$:
$\Delta_{t}=G_{t}^{2}$ $\leq\frac{\mathcal{A}^{2}(1-p)^{2}}{C^{2}}\cdot\frac{(\ln t)^{2}}{t^{2(1-p)}}$
$=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right).$
@@
$=\Delta_{t}-\frac{\sqrt{2\mu}}{t^{3/4}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
By Lemma E.2, let $G_{t}=\sqrt{\Delta_{t}}$, $C=\sqrt{2\mu}/2$ and $p=3/4$.
$G_{t+1}\leq G_{t}-\frac{C}{t^{3/4}}+\frac{\Gamma_{t}}{2G_{t}}.$
Setting $p=3/4$, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{8C}$, there exists a time step $T_{A}$ such that for all $t>T_{A}$, we have $G_{t}\leq A\cdot\frac{\ln t}{t^{1/4}}$. By choosing $A=\frac{\mathcal{A}}{4C}$, we satisfy the condition $A>\frac{\mathcal{A}}{8C}$, which ensures this bound holds. Thus, we obtain the convergence rate for $\Delta_{t}$:
$\Delta_{t}\leq\frac{\mathcal{A}^{2}(1-3/4)^{2}}{C^{2}}\cdot\frac{(\ln t)^{2}}{t^{2(1-3/4)}}=\frac{\mathcal{A}^{2}}{8\mu}\cdot\frac{(\ln t)^{2}}{t^{1/2}}$
We can now summarize the results for both scenarios. In either case ($\gamma=0$ or $\gamma\neq 0$), the analysis yields the same asymptotic convergence rate. For any number of iterations $T>T_{0}=2e^{4}$, the analysis leads to the same asymptotic upper bound:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{\mathcal{A}^{2}}{8\mu}\cdot\frac{(\ln T)^{2}}{T^{1/2}}=\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right).$
The specific definition of $\mathcal{A}$ is as follows: (i) Case 1 ($\gamma=0$): $\mathcal{A}=6L^{-1}\sigma^{2}+8\sqrt{2}Ln+2Ln+L$. (ii) Case 2 ($\gamma\neq 0$): $\mathcal{A}=14L^{-1}\sigma^{2}+16\sqrt{2}Ln+2Ln+L$.
@@
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}.$
@@
$+\underbrace{\frac{L\eta_{t}^{3/2}}{2}+\left(\frac{\epsilon}{2}+\frac{\sqrt{\eta_{t}}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}}_{\Gamma_{t}},$
@@
$\Gamma_{t}=\frac{\sqrt{\eta_{t}}}{L}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+(L/2+Ln)\eta_{t}^{3/2},$
@@
$\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}.$
@@
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{2/3}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
This recurrence is in the form required by Lemma E.2. Let $G_{t}=\sqrt{\Delta_{t}}$, $C=\sqrt{2\mu}/2$ and $p=2/3$. According to Lemma E.2, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{6C}$, there exists a time $T_{A}$ such that for all $t>T_{A}$, $G_{t}\leq A\cdot\frac{\ln t}{t^{1/3}}$. We choose a specific constant $A=\frac{\mathcal{A}}{3C}$, where $\mathcal{A}:=A_{1}+A_{2}=20L^{-1}\sigma^{2}+34Ln+L$. Squaring both sides of the inequality for $G_{t}$ establishes the convergence rate for $\Delta_{t}=G_{t}^{2}$:
$\Delta_{t}\leq A^{2}\cdot\frac{(\ln t)^{2}}{t^{2/3}}=\frac{\mathcal{A}^{2}}{9C^{2}}\frac{(\ln t)^{2}}{t^{2/3}}=\frac{2\mathcal{A}^{2}}{9\mu}\frac{(\ln t)^{2}}{t^{2/3}}$
Therefore, for any number of iterations $T>T_{0}=2e^{3}$, the expected suboptimality has the following asymptotic bound:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{2\mathcal{A}^{2}}{9\mu}\cdot\frac{(\ln T)^{2}}{T^{2/3}}=\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right).$